NOAA SATOPS Morning Report: Monday, February 11, 2013


SATOPS Morning Report: Monday, February 11, 2013

Customers or users who want to call to ask about problems or report outages can use the following contacts:
- For questions about data quality (noisy data, dropouts, schedules), GOES or POES products, or status of spacecraft can call the SAB Shift Supervisor at 301-763-8444
- To report an outage or document a problem, call the ESPC Operations Crew Lead at 301-457-5218. All LRIT data users are encouraged to report problems to: LRIT@noaa.gov

DOY 042

Critical Weather Day ended Feb. 10, 0000Z (declared for NE Snowstorm).

On Friday, February 8, 2013, at approximately 11:40 am local time, the NSOF facility lost commercial power; the operations floor and its systems converted over to generator power with no adverse impact to
operational equipment. Commercial power was restored to the NSOF facility @ approximately 12:55 pm local time.

GOES
GOES 12 IMAGES AND SOUNDING CANCELLATIONS FOR KEEP OUT ZONE:
0315, 0327, 0345, 0357, 0415, 0427, 0445Z.

All other GOES operations were nominal over the past 72 hours.
No anomalies affecting product delivery via GVAR or s/c health & safety.
For GOES status and other information please see:
http://www.oso.noaa.gov/goes

POES
All POES operations were nominal over the past 72 hours.
No anomalies affecting product delivery or s/c health & safety.
For a detailed status and description of POES see:
http://www.oso.noaa.gov/poes

DMSP
Day 040: Due to a ground system configuration issue at LION-C, Ops 54 rev. 48050 06:18Z, had a SMOOTH CAT 3 Data Loss (s/c sec: 7218/ Pb sec: 245) and TF CAT 2 Data Loss (s/c sec: 1168/ Pb sec: 292).

Day 041: Two passes (OPS 53, rev 17102 and OPS 47, rev 92345) were scheduled on the same FBKS antenna, using different links, which yielded a conflict between real-time operations on OPS 53 & prepass operations on OPS 47. CAT 1 data loss on OPS 53 (51 PB secs/ 1500 SC secs).

Day 042: Problem with OPS 53 rev 17107; AFWA reports numerous issues with ingesting GIGS data from OPS 53. Cat 2 Data Loss on OPS 53 (122 SC seconds/ 61 PB seconds). Other bad GIGS data recovered at other sites.

All other DMSP operations were nominal over the past 72 hours.
No anomalies affecting product delivery or s/c health & safety.
For more DMSP Information please see:
http://dmsp.ngdc.noaa.gov/dmsp.html

JASON-2
Due to an AMR instrument issue, the Jason-2 OGDR Products distributed Feb. 8, 2013 were degradedfrom sensing time 17:37:36Z to 22:46:50Z (5 hr 9 min 14 sec):
JA2_OPN_2PdS169_176_20130208_171438_20130208_191120.nc
JA2_OPN_2PdS169_178_20130208_191119_20130208_205102.nc
JA2_OPN_2PdS169_179_20130208_205101_20130208_224619.nc
JA2_OPN_2PdS169_181_20130208_224618_20130209_004315.nc (first 11 secs)

All other JASON-2 operations were nominal over the past 72 hours.
No anomalies affecting product delivery or s/c health & safety.

SNPP
All SNPP operations were nominal over the past 72 hours.
No anomalies affecting product delivery or s/c health & safety.

ESPC Reported data outage and delays
ESPC did not received HRPT data from Monterey February 8, 2013 from 0245z to 2040z. Monterey had a problem with their network switch.

AQUA AIRS data outage from February 8, 2013 21:32Z until February 10, 2013 01:15Z.

Questions or comments to Ron.Mahmot@noaa.gov

Please follow SpaceRef on Twitter and Like us on Facebook.